OpenStudy (anonymous):

F(4)=6 f^-1 (6)=4 F'(4)=8 F'(6)=10 compute the derivative of the inverse function F

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

the slope at (4,6) is 8

OpenStudy (amistre64):

cant see it at the moment..

OpenStudy (amistre64):

F(4)=6 F^-1(6)=4 F'(4)=8 .... F'(6)=10 ...

OpenStudy (amistre64):

1/8 and 1/10 is what i wanna say

OpenStudy (amistre64):

an inverse flips over the y=x line

OpenStudy (anonymous):

I actually just got the same answer

OpenStudy (amistre64):

(dx/dy) = 1/(dx/dy) is what i thought :)

OpenStudy (amistre64):

well, something like that lol

OpenStudy (anonymous):

lol, thanks

OpenStudy (amistre64):

yw :)

OpenStudy (amistre64):

for example: x^2 derives to 2x; at(3,9) the slope is 6 inverse it: sqrt(y) derives to 1/2sqrt(y) at(9,3) = 1/2sqrt(9) = 1/6

OpenStudy (anonymous):

No need to explain it lol. I understand the concept of inverse-ing functions and their derivatives.

OpenStudy (anonymous):

Thanks anyways

OpenStudy (amistre64):

that was more for me then :) enjoy ;)